# Canonical Representation and Force-Based Pretraining of 3D Tactile for Dexterous Visuo-Tactile Policy Learning

### v1

@@ V-F Role of Spatial Information and Force Information
We conducted an ablation study to validate the use of spatial and force information in our policy. In the spatial ablation, we kept the observation of each tactile sensor’s 6D pose fixed at its initial state during manipulation, while in the force ablation, we set the observation of all tactile forces to zero. In both cases, the robot failed to flip the bottle cap at all. In the spatial ablation, once the robot reached the object and attempted to grasp it, the thumb oscillated randomly, preventing further manipulation. In the force ablation, although the robot reached the object and attempted to grasp it, it consistently failed due to an unstable grasp or continuous adjustments. These results demonstrate that our policy leverages spatial information for forming gross hand poses and force information for more fine-grained adjustments.
V-G Generlization
To validate the generalization of our method, we tested the policy with four unseen objects varying in color, geometry, and dynamics, with each object being tested twice for the open box and flip tasks. As shown in Fig. 5, our policy successfully opens the box 5 times out of 8 tries. For one failure of the open box task, although the hand opened the box to a certain degree that generally won’t fall down, it fell down due to completely different friction properties of the box. For the flip task, the policy succeeded 6 times out of 8 tries, demonstrating the generalization ability of our method.

### v2

@@ V-F Role of Spatial Information and Force Information
We conducted an ablation study to validate the use of spatial and force information in our policy. In the spatial ablation, the tactile sensor’s 6D pose was fixed at the initial state, while in the force ablation, all tactile forces were set to zero. In both cases, the robot failed to flip the cap at all. In the spatial ablation, once the robot reached the object and attempted to grasp it, the thumb oscillated randomly, preventing further manipulation. In the force ablation, although the robot reached the object and attempted to grasp it, it consistently failed due to an unstable grasp or continuous adjustments. These results demonstrate that our policy leverages spatial information for forming gross hand poses and force information for more fine-grained adjustments.
V-G Generlization to Unseen Objects
To validate the generalization of our method, we tested the policy with four unseen objects exhibiting diverse color, geometry, and dynamics, with each object being tested twice for the open box and flip tasks. As shown in Fig. 5, our policy successfully opens the box 5 times out of 8 tries. For one failure of the open box task, although the hand opened the box to a certain degree that generally won’t fall down, it fell down due to completely different friction properties of the box. For the flip task, the policy succeeded 6 times out of 8 tries, demonstrating the generalization ability of our method.
@@ VI CONCLUSIONS
ACKNOWLEDGMENT
This work is supported by the National Youth Talent Support Program (8200800081) and National Natural Science Foundation of China (No. 62376006).